OpenStudy (anonymous):

Solve and show work h-48+17h=42 h=?

OpenStudy (anonymous):

solve for h first combine the like terms on the left side of the equation

OpenStudy (goldphenoix):

\[\large \large h-48+17h=42\] You want to combine like term: 17h, and h are like term. This will give you 18h. Now your equation should look like: -48+18h = 42 Now you want to get rid of -48 to solve for h. So: -48+18h = 42 +48 +48 ------------------ 18h = 90 Last step is to divide!!!: 18h = 90 Divide 18 to get 5. So h = 5! Hope this help!

OpenStudy (anonymous):

h=5